# Time-bin entanglement in the deterministic generation of linear photonic cluster states

### v2

@@ Abstract
We investigate strategies for the efficient deterministic creation of trains of time-bin entangled photons using an individual quantum emitter described by a $\Lambda$-type electronic system. We explicitly demonstrate generation of high-quality linear cluster states of substantial length in our full microscopic numerical simulations. The underlying scheme is based on the manipulation of ground state coherences through precise optical driving. One important finding is that the most easily accessible quality metrics, the achievable rotation fidelities, fall short in assessing the actual quantum correlations of the emitted photons in the face of losses. To address this, we explicitly calculate stabilizer generator expectation values as a superior gauge for the quantum properties of the many-photon state. Our results illustrate that with controlled minimization of losses and realistic system parameters for quantum-dot type systems, useful linear cluster states of significant lengths can be generated, showcasing promise of scalability for quantum information processing endeavors.
@@
Quantum information theory and processing are poised to fundamentally transform computational paradigms, surpassing the limitations of classical systems [1]. Central to this quantum leap is the development of reliable and deterministic photon emitters [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], which are essential for the encoding, transmission, and manipulation of quantum information in measurement-based quantum computing [18, 19] and quantum communication [20, 21, 22, 23, 24, 25, 26, 27]. Multi-photon states with significant quantum correlations or entanglement are a critical ingredient for numerous quantum protocols [22, 28, 29] and the quest for their robust generation forms a challenging field of research [30, 31]. Besides the inherently limited polarization [32, 28] and other degrees of freedom [33, 34, 35, 36], quantum correlations can be encoded in the form of time-bin entanglement [37, 38, 39, 40, 41] of photonic qubits [38, 42], which holds great potential for applications in quantum teleportation and quantum key distribution [43, 39, 44].
@@ I $\Lambda$-Type Emitter Structure
The design of the emitter structure is pivotal, as it must exhibit a $\Lambda$-shaped energy configuration of the underlying electronic states. Such a configuration enables us to achieve dual objectives: firstly, to excite individual electronic states, leading to the emission of photons via radiative decay; and secondly, to infuse the emitted photons with the desired quantum phase information, a direct consequence of the coherences between the states within the $\Lambda$ system, which is schematically depicted in Fig. 1. The capability to instill quantum coherence is essential for the generation of photon entanglement. The temporal entanglement of photonic qubits, as produced by this structure, plays a foundational role in a wide array of quantum protocols, and thus constitutes a key mechanism for advancing the field of quantum information science. For example, specific types of quantum dot molecules are known to exhibit such energy structure, making them a prime solid-state candidate for realizing the $\Lambda$-type configurations required.
@@
Figure 2: Energy-level structure and emission scheme. a) The $\Lambda$-System depicted, which may for example be realized in quantum dot molecule systems, features two ground states which are separated by a finite energy splitting $\Delta_{\text{GS}}=E_{G_{2}}-E_{G_{1}}$. Both ground states $G_{1}$ and $G_{2}$ couple to excited states, which can emit a single photon through radiative decay with rate $\gamma_{\text{Cyc}}$; for simplicity only the first excited state is used for photon emission here. Consequently, the second excited state is depicted in a faded manner in the illustration. Both excited states can be driven using a laser pulse $\Omega_{\text{Cyc}}$. The transition states are driven by two additional laser pulses $\Omega_{1,2}$ and are split energetically by $\epsilon$. The desired rotations in the coherent superposition of the two ground states can be achieved by adjusting the transition pulse detunings using $\delta>0$. b) Emission sequence simulated for six photons as an example; each photon is emitted into a time bin, pairs of subsequent early and late time bins form a photonic qubit. Quantum correlations are defined between the states of these qubit. Pulses are hyperbolic sechant shaped, with lengths of 16ps for rotation and 5ps for extraction pulses. Radiative decay is present for the cycling transitions in the scenario shown here, losses for other states and transitions are included further below as discussed in the text. Inset: Bloch sphere illustrating rotations in the x, y plane and along the z axis.
@@
The specific $\Lambda$-type system investigated here is characterized by its two distinct ground states, the excited states and the transition and loss states. Regardless of the physical implementation, the ground states will exhibit some degree of spectral splitting $\Delta_{\text{GS}}$. We will call the ground state that is spectrally closer to the transition states first ground state, and the corresponding other second ground state. The ground state splitting can be achieved e.g. by using stacked quantum dots. Each of the ground states can be individually driven by differently polarized laser pulses $\Omega_{\text{Cyc}}$, offering high excitation fidelity of each specific cycling transition. The excited states will exhibit radiative decay; here we assume a radiative lifetime of $0.51/\gamma_{\text{Cyc}}=$ 0.5ns, which results in a full recombination of the excited states within about $3$ns. This is depicted in Fig. 2 b) and lies well within the realistic range for commonly manufactured quantum dots [57, 58, 36, 35]. For the sake of simplicity in our analysis, we only consider photon emission from the excited state associated with the first ground state, even though our numerical model also includes the potential for emissions linked to the second ground state’s excited state. Both ground states couple to several transition states. In this case, the transition states are limited to a single target transition and a single unwanted loss transition, slightly reducing rotation fidelities. Moreover, both ground states couple to a superposition of available transition states. These transitions are driven by the optical pulses $\Omega_{1/2}$, targeting both the target and the unwanted state. We set the excited state energy as well as the transition state energy to a fixed energy $\hbar\omega=1$eV, which is well within the values for commonly used QD materials [58, 36, 5]. We note that the system performs analogously also for different energies. We further adopt viable system parameters from previous work [50], setting the target-unwanted splitting to $\epsilon=500\upmu$eV and fixing the ground state splitting to $\Delta_{\text{GS}}=200\upmu$eV. Setting both values to zero reduces the transition system to a simple degenerate $\Lambda$ system, avoiding the need to phase tune the rotation pulses. By fixing parameters such as the excited and transition state energies, along with specific energy splittings, we intentionally limit the variables in our study. This approach streamlines our focus on essential factors that significantly impact the system’s ability to generate entangled streams of photons (avoiding a full optimization of the system, which could be a part of future work). For the rotation pulses, we use hyperbolic secant shaped envelopes defined by
@@
$\Omega_{i}(t)=\alpha\sigma\text{sech}\left[\sigma(t-t_{i})\right]\cdot\text{exp}\left[-i(\omega_{i}(t-t_{i})+\vartheta_{i})\right]~{},$ (1)
@@
with pulse amplitude $\alpha$, pulse bandwidth $\sigma$, center $t_{i}$ and pulse frequency $\omega_{i}$. Through the bandwidth, we set the pulse length to a fixed $\tau_{i}=\frac{1}{\sigma}=16$ps for the transition pulses and $\tau_{\text{Cyc}}=5$ps for the excitation pulses. The pulse width translates into a pulse bandwidth of $\approx 41\upmu$$/\hbar$, which, in previous work [43], proved to be around the ideal bandwidth for the given energy configuration. To achieve a desired rotation of the ground state superposition by angle $\Theta$, we set the pulse amplitude of the total driving force of the system to $\alpha_{1}^{2}+\alpha_{2}^{2}=1$ and use a finite non zero detuning of the pulse energy. Prior work has shown that equal coupling to both states is most effective in achieving high fidelity rotations of the ground state coherences [43]. Therefore, we limit our investigation to cases of equal coupling where both pulses couple equally to both of the transition states. This means, that both the unwanted and the target state pulses use the same amplitude and phase, while the phases between pulses driving transitions for different ground states can differ. However, we note that a generalization to arbitrary coupling rates is easily achievable and merely increases the size of the parameter space to be analyzed.
@@
We simulate the temporal dynamics of the system using the von Neuman equation $\frac{\partial d\rho}{\partial t}=-\frac{i}{\hbar}[H,\rho]+\frac{\gamma_{O}}{2}\sum_{\hat{O}}2\hat{O}\rho\hat{O}^{\dagger}-\hat{O}^{\dagger}\hat{O}\rho-\rho\hat{O}^{\dagger}\hat{O}$ with Lindblad-type losses, where $\hat{O}$ denotes any of the available system state operators and $\gamma_{O}$ is the corresponding loss rate. The density matrix $\rho$ in matrix representation is propagated in time using explicit numerical integration methods. We distinguish between the following decays or loss channels: the radiative decay of the cycling transition ($\gamma_{\text{Cyc}}$), the decay of both the target and unwanted state population ($\gamma^{R}_{T,U}$), and the spin dephasing rate of the ground states ($\gamma^{D}_{T,U}$). Unless specified otherwise, we assume that both $\gamma^{D}_{T,U}$ and $\gamma^{R}_{T,U}$ are set to zero.
@@
Figure 3: Sensitivity to detuning of rotation pulses. Panels a-c) show Bloch-sphere representations of the rotations annotated in d) from left to right. The blue arrow indicates the initial state, the orange arrow indicates the final state. The spheres show rotations for the lossless case with normalized state vectors. Panel d) shows expectation values for the first and second ground state population for varying pulse detuning of the rotation pulses; the pulse amplitudes are fixed to $\Omega_{\text{1,2}}=\frac{\pi}{\sqrt{2}}$. For a detuning of zero, both pulses are resonant with their respective transitions. The inset e) depicts a zoom on the $\pi$ rotation. Dotted lines show rotations with exaggerated relaxation of the target and unwanted states ($\gamma^{R}_{T,U}=2\gamma_{\text{Cyc}}$, green dotted line). Dashed lines show rotations with exaggerated spin dephasing of the ground states ($\gamma^{D}_{T,U}=\gamma_{\text{Cyc}}$, red dashed line).
@@ II Time-Bin Emission
  • Start or prepare the system in $|\Psi\rangle=\frac{1}{\sqrt{2}}\left(|G_{1}\rangle+|G_{2}\rangle\right)$
@@
Steps 1-4 are repeated N times to create an N qubit linear cluster states. The final step, involving the removal of the emitter from the cluster state, is not always necessary and depends on the specific desired state [53]. The photon emission occurs through spontaneous radiative emission from the corresponding excited state. The x-y angle $\vartheta$ is tunable by selecting appropriate pulse phases, while $\Theta$ can be selected varying the pulse amplitudes or detunings. We consistently select a bin length that accurately encompasses the emitted photons, in accordance with $\gamma_{\text{Cyc}}$. For a radiative decay rate of $\gamma_{\text{Cyc}}=1.2\upmu$eV ($\approx 2$GHz), we find that a time bin length of $T=4$ns suits the emission scheme well as it incorporates both the emission process as well as a time buffer to apply the rotation pulses. Additionally, we choose rotation pulses of the highest fidelity, ensuring that the pulse bandwidth is appropriately matched to the selected splittings. In Fig. 2 b), the emission of up to six photons is illustrated; it is important to note though that the system or protocol are not constrained to this particular number of photons.
@@
We observe that the electronic coherences disappear immediately upon the application of the $\pi$-extraction pulse through the cycling transition, provided the system is in an equal superposition of both ground states. Further $\Theta=\pi$ or $\Theta=\pi/2$ rotations through the $\Lambda$-system are then not visible withing the electronic coherences and, in an ideal case, leave the state population constant. For this reason, these pulses are referred to as transitionless [43]. This phenomenon is by design, as it facilitates the transfer of coherences from the electronic states onto the emitted photons.
@@ III Stabilizer Generator Expectation Values
For a full quantum state tomography or detailed analysis of photon entanglement in a given N photon quantum state, one might consider computing the N-photon density matrix directly. From here, protocol fidelities can be calculated [68]. However, in the microscopic simulations, this approach becomes computationally intractable, typically already for $N>3$, due to the exponential increase in space dimensionality and requirement for the explicit evaluation of N-th order photon correlation functions $\mathcal{G}^{(N)}$. To keep the computational effort manageable, here we utilize stabilizer generator expectation values [62] as our primary metric for evaluating quantum correlations in the generated N photon quantum states of interest. These values allow a localized investigation of entanglement by probing correlations in a segmented fashion, thereby significantly reducing computational demands. This is accomplished by evaluating correlation functions $\mathcal{G}^{(3)}$ only up to third-order, resulting in a tractable and efficient means to estimate the quantum entanglement present in the system without the need to reconstruct the full N-photon state. This effectively local approach not only simplifies the computational process but also delivers a detailed assessment of the entanglement properties. Stabilizer generator expectation values offer these capabilites to serve as benchmarks for these states because by definition, the graph states are composed of a superposition of stabilizer expressions [62, 69]. Unlike readily accessible metrics for generation protocol quality, such as rotation fidelities, stabilizer generator expectation values that are directly derived from the N-photon quantum state offer an accurate reflection of the system’s entanglement characteristics [8], even in lossy environments. To construct the stabilizer generator expectation values within the early and late basis specifically for a 1D linear cluster state, we explicitly calculate both second and third order photon correlation functions:
@@
where time arguments $t_{i}$ are shifted by the time bin length $T$, as indicated by the $E$ (early) and $L$ (late) subsripts of the corresponding correlation functions (cf. Appendix A). The multi-time correlation functions are calculated by using the von-Neuman equation again on a modified density matrix using the quantum regression theorem [35, 70]. The operator $a^{(\dagger)}$ is the annihilation (creation) operator corresponding to the first ground- and excited state. We refer to the transition between these states as the cycling transition. We then reduce the correlation functions by integrating over all time dimensions such that
@@
The operator $\hat{\Phi}=\cos(\varphi)X+i\sin(\varphi)Y$ defines the phase $\varphi$ of the stabilizer generator expectation values which is linearly tunable by adjusting the phase $\vartheta$ of the rotation pulse. The numerator is normalized with
@@
These expressions are specific to 1D linear cluster states, grow linearly with the temporal photon entanglement [39] and can be used to calculate a lower bound for the generated entanglement. In the ideal case, the stabilizer generator expectation values approach unity. The necessary configuration for the local bins to calculate any of these stabilizer generator expectation values adheres to the sequence
@@
where $t_{0}$ marks the commencement of the first bin associated with the expression, and its positioning is flexible within the time frame of the simulation to calculate later generator expectation values. Further details are given in Appendix A.
@@
We now explore how rotation fidelities directly influence the quantum characteristics of the generated N photon states, specified by the stabilizer generator expectation values. Given that it is always possible to identify a pulse phase optimizing both characteristics simultaneously, we select an arbitrary phase for our analysis, focusing solely on the magnitude of these properties. Initially, we introduce additional radiative loss exclusively to the transition states, previously set to zero, and adjust it from zero to $2.5\upmu$eV. The radiative loss rate for the cycling transition is kept constant. In Fig. 5, we observe that increased radiative losses in the transition state correspond to a direct decrease in the magnitudes of the stabilizers, exhibiting an almost linear relationship for the losses shown. Subsequently, we apply spin dephasing to the ground state coherences, again with rates ranging from zero to $2.5\upmu$eV. We note that these values lead to dephasing times significantly exceeding the typical achievable values for spin qubits [71] and provide a solid estimation for the lower bound for the achievable entanglement. In this scenario, we notice a rapid exponential decline in the stabilizer values even for relatively small dephasing rates. It is important to note that this detrimental effect of the ground state dephasing is not apparent in the temporal dynamics of the system, which for small loss values closely resembles the zero-dephasing behavior depicted in Fig. 2 b). It is only for losses exceeding $3\upmu$eV that a noticeable difference also becomes visible for the system dynamics.
@@
Figure 6: Local stabilizer generator expectation values and cluster-state length limitation. a) Magnitudes of the first 15 $\langle Z\Phi Z\rangle$ expectation values. The lossless case (blue line) is compared to either significant decay of the target and unwanted states (orange line) with $\gamma^{R}_{T,U}=0.5\cdot\gamma_{\text{Cyc}}\equiv 1$GHz, which corresponds to a lifetime of the transition states of $1$ns, or moderate spin dephasing of the ground states (green line) with $\gamma^{D}_{T,U}=0.0375\cdot\gamma_{\text{Cyc}}\equiv 75$MHz, which corresponds to a spin lifetime greater than $13$ns. b) Lower bound for the possible length of the linear cluster state according to Eq. 13, assuming all stabilizer expectation values are equal.
@@
It becomes evident that, although the rotation fidelity diminishes somewhat with increasing losses, the nonlinear decay of quantum correlations with increasing spin dephasing does only emerge in a detailed analysis of the quantum stabilizer generators. We find that the selection of an optimal bin length is also influenced by the system’s loss parameters (not shown), with systems experiencing higher levels of loss potentially benefiting from shorter time bins, which helps to mitigate the adverse effects of dephasing over extended periods of time, optimizing the overall performance of the system under such conditions.

### v3

@@ Abstract
We theoretically investigate strategies for the deterministic creation of trains of time-bin entangled photons using an individual quantum emitter described by a $\Lambda$-type electronic system. We explicitly demonstrate the theoretical generation of linear cluster states with substantial numbers of entangled photonic qubits in full microscopic numerical simulations. The underlying scheme is based on the manipulation of ground state coherences through precise optical driving. One important finding is that the most easily accessible quality metrics, the achievable rotation fidelities, fall short in assessing the actual quantum correlations of the emitted photons in the face of losses. To address this, we explicitly calculate stabilizer generator expectation values as a superior gauge for the quantum properties of the generated many-photon state. With widespread applicability also to other emitter and excitation-emission schemes, our work lays the conceptual foundations for an in-depth practical analysis of time-bin entanglement based on full numerical simulations with predictive capabilities for realistic systems and setups including losses and imperfections. The specific results shown in the present work illustrate that with controlled minimization of losses and realistic system parameters for quantum-dot type systems, useful linear cluster states of significant lengths can be generated in the calculations, discussing the possibility of scalability for quantum information processing endeavors.
@@
Quantum information theory and processing are poised to fundamentally transform computational paradigms, surpassing the limitations of classical systems [1]. Central to this quantum leap is the development of reliable and deterministic photon emitters [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], which are essential for the encoding, transmission, and manipulation of quantum information in measurement-based quantum computing [18, 19] and quantum communication [20, 21, 22, 23, 24, 25, 26, 27]. Multi-photon states with significant quantum correlations or entanglement are a critical ingredient for numerous quantum protocols [22, 28, 29] and the quest for their robust generation forms a challenging field of research [30, 31]. Besides polarization [32, 28] and other degrees of freedom [33, 34, 35, 36], quantum correlations can be encoded in the form of time-bin entanglement [37, 38, 39, 40, 41] of photonic qubits [38, 42], which holds great potential for applications in quantum teleportation and quantum key distribution [43, 39, 44].
@@ I $\Lambda$-Type Emitter Structure
The design of the emitter structure is pivotal, as it must exhibit a $\Lambda$-shaped energy configuration of the underlying electronic states. Such a configuration enables us to achieve dual objectives: firstly, to excite individual electronic states, leading to the emission of photons via radiative decay; and secondly, to infuse the emitted photons with the desired quantum phase information, a direct consequence of the coherences between the states within the $\Lambda$ system, which is schematically depicted in Fig. 1. The capability to instill coherence of the quantum state is essential for the generation of photon entanglement. The temporal entanglement of photonic qubits, as produced by this structure, plays a foundational role in a wide array of quantum protocols, and thus constitutes a key mechanism for advancing the field of quantum information science. For example, specific types of quantum dot molecules are known to exhibit such energy structure, making them a prime solid-state candidate for realizing the $\Lambda$-type configurations required.
@@
Figure 2: Energy-level structure and emission scheme. a) The $\Lambda$ system depicted, which may for example be realized in quantum dot molecule systems, features two ground states which are separated by a finite energy splitting $\Delta_{\text{GS}}=E_{G_{2}}-E_{G_{1}}$. Both ground states $G_{1}$ and $G_{2}$ couple to excited states, which can emit a single photon through radiative decay with rate $\gamma_{\text{Cyc}}$; for simplicity only the first excited state is used for photon emission here. Consequently, the second excited state is depicted in a faded manner in the illustration. Both excited states can be driven using a laser pulse $\Omega_{\text{Cyc}}$. The transition states are driven by two additional laser pulses $\Omega_{1,2}$ and are split energetically by $\epsilon$. The desired rotations in the coherent superposition of the two ground states can be achieved by adjusting the transition pulse detunings using $\delta>0$. b) Emission sequence simulated for a combined total of three photons as an example; each photon is emitted into a time bin, pairs of subsequent early and late time bins form a photonic qubit. Quantum correlations are defined between the states of these qubit. Pulses are hyperbolic sechant shaped, with lengths of 16ps for rotation and 5ps for extraction pulses. Radiative decay is present for the cycling transitions in the scenario shown here, losses for other states and transitions are included further below as discussed in the text. Inset: Bloch sphere illustrating rotations in the x, y plane and along the z axis.
@@
The specific $\Lambda$-type system investigated here is characterized by its two distinct ground states, the excited states and the transition and loss states. Regardless of the physical implementation, the ground states will exhibit some degree of spectral splitting $\Delta_{\text{GS}}$. We will call the ground state that is spectrally closer to the transition states first ground state, and the corresponding other second ground state. The ground state splitting can be achieved e.g. by using stacked quantum dots. Small alignment mismatches will then result in mixing of the states, enabling rotations through the transition states. Each of the ground state transitions can be individually driven by differently polarized laser pulses $\Omega_{\text{Cyc}}$, offering high excitation fidelity of each specific cycling transition. The excited states will exhibit radiative decay; here we assume a radiative lifetime of $0.51/\gamma_{\text{Cyc}}=$ 0.5ns, which results in an almost full recombination of the excited states within about $3$ns. This is depicted in Fig. 2 b) and lies well within the realistic range for commonly manufactured quantum dots [57, 58, 36, 35]. For the sake of simplicity in our analysis, we only consider photon emission from the excited state associated with the first ground state, even though our numerical model also includes the potential for emissions linked to the second ground state’s excited state. Both ground states couple to several transition states. In this case, the transition states are limited to a single target transition and a single unwanted transition, slightly reducing rotation fidelities. Moreover, both ground states couple to a superposition of available transition states. These transitions are driven by the optical pulses $\Omega_{1/2}$, targeting both the target and the unwanted state. We set the excited state energy as well as the transition state energy to a fixed energy $\hbar\omega=1$eV, which is well within the values for commonly used QD materials [58, 36, 5]. We note that the system performs analogously also for different energies. We further adopt viable system parameters from previous work [50], setting the target-unwanted splitting to $\epsilon=500\upmu$eV and fixing the ground state splitting to $\Delta_{\text{GS}}=200\upmu$eV. For definition of the system in form of its Hamiltonian see Appendix A. Setting both values to zero reduces the transition system to a simple degenerate $\Lambda$ system, avoiding the need to phase tune the rotation pulses. By fixing parameters such as the excited and transition state energies, along with specific energy splittings, we intentionally limit the variables in our study. This approach streamlines our focus on essential factors that significantly impact the system’s ability to generate entangled streams of photons (avoiding a full optimization of the system, which could be a part of future work). For the rotation pulses, we use hyperbolic secant shaped envelopes defined by
@@
$\Omega(t)=\alpha\sigma\text{sech}\left[\sigma(t-t_{0})\right]\cdot\text{exp}\left[-i(\omega(t-t_{0})+\vartheta)\right]~{},$ (1)
@@
with pulse amplitude $\alpha$, pulse bandwidth $\sigma$, center $t_{0}$ and pulse frequency $\omega$. Through the bandwidth, we set the pulse length to a fixed $\tau_{\text{Transition}}=\frac{1}{\sigma}=16$ps for the transition pulses and $\tau_{\text{Cyc}}=5$ps for the excitation pulses. The pulse width translates into a pulse bandwidth of $\approx 41\upmu$$/\hbar$, which, in previous work [43], proved to be around the ideal bandwidth for the given energy configuration. To achieve a desired rotation of the ground state superposition by angle $\Theta$, we set the pulse amplitude of the total driving force of the system to $\alpha_{1}^{2}+\alpha_{2}^{2}=1$ and use a finite non zero detuning of the pulse energy. Prior work has shown that equal coupling to both states is most effective in achieving high fidelity rotations of the ground state coherences [43]. Therefore, we limit our investigation to cases of equal coupling where both pulses couple equally to both of the transition states. This means, that both the unwanted and the target state pulses use the same amplitude and phase, while the phases between pulses driving transitions for different ground states can differ. However, we note that a generalization to arbitrary coupling rates is easily achievable and merely increases the size of the parameter space to be analyzed.
@@
We simulate the temporal dynamics of the system using the von Neuman equation $\frac{\partial\rho}{\partial t}=-\frac{i}{\hbar}[H,\rho]+\sum_{\hat{O}}\frac{\gamma_{O}}{2}\left(2\hat{O}\rho\hat{O}^{\dagger}-\hat{O}^{\dagger}\hat{O}\rho-\rho\hat{O}^{\dagger}\hat{O}\right)$ with Lindblad-type losses, where $\hat{O}$ denotes any of the available system state operators and $\gamma_{O}$ is the corresponding loss rate. The density matrix $\rho$ in matrix representation is propagated in time using explicit numerical integration methods and using the Hamiltonian in Eq. 16. We distinguish between the following decays or loss channels: the radiative decay of the cycling transition ($\gamma_{\text{Cyc}}$), the decay of both the target and unwanted state population ($\gamma^{\text{R}}_{\text{T,U}}$), and the spin dephasing rate of the ground states ($\gamma^{\text{D}}_{\text{T,U}}$). Unless specified otherwise, we assume that both $\gamma^{\text{D}}_{\text{T,U}}$ and $\gamma^{\text{R}}_{\text{T,U}}$ are set to zero.
@@
Figure 3: Sensitivity to detuning of rotation pulses. Panels a-c) show Bloch-sphere representations of the rotations annotated in d) from left to right. The blue arrow indicates the initial state, the orange arrow indicates the final state. The spheres show rotations for the lossless case with normalized state vectors. Panel d) shows expectation values for the first and second ground state population for varying pulse detuning of the rotation pulses; the pulse amplitudes are fixed to $\Omega_{\text{1,2}}=\frac{\pi}{\sqrt{2}}$. For a detuning of zero, both pulses are resonant with their respective transitions. The inset e) depicts a zoom on the $\pi$ rotation. Dotted lines show rotations with exaggerated relaxation of the target and unwanted states ($\gamma^{\text{R}}_{\text{T,U}}=2\gamma_{\text{Cyc}}$, green dotted line). Dashed lines show rotations with exaggerated spin dephasing of the ground states ($\gamma^{\text{D}}_{\text{T,U}}=\gamma_{\text{Cyc}}$, red dashed line).
@@ II Time-Bin Emission
  • Start or prepare the system in $|\Psi\rangle=\frac{1}{\sqrt{2}}\left(|\text{G}_{1}\rangle+|\text{G}_{2}\rangle\right)$
@@
Steps 1-4 are repeated N times to create an N qubit linear cluster states. The final step, involving the removal of the emitter from the cluster state, is not always necessary and depends on the specific desired state [53]. The photon emission occurs through spontaneous radiative emission from the corresponding excited state. The x-y angle $\vartheta$ is tunable by selecting appropriate pulse phases, while $\Theta$ can be selected varying the pulse amplitudes or detunings. We consistently select a bin length that accurately encompasses the emitted photons, in accordance with $\gamma_{\text{Cyc}}$. For a radiative decay rate of $\gamma_{\text{Cyc}}=1.2\upmu$eV ($\approx 2$GHz), we find that a time bin length of $T=4$ns suits the emission scheme well as it incorporates both the emission process as well as a time buffer to apply the rotation pulses. Additionally, we choose rotation pulses of the highest fidelity, ensuring that the pulse bandwidth is appropriately matched to the selected splittings. In Fig. 2 b), the emission of up to three photons is illustrated; it is important to note though that the system or protocol are not constrained to this particular number of photons.
@@
We observe that the electronic coherences disappear immediately upon the application of the $\pi$-extraction pulse through the cycling transition, provided the system is in an equal superposition of both ground states. Further $\Theta=\pi$ or $\Theta=\pi/2$ rotations through the $\Lambda$-system are then not visible withing the electronic coherences and, in an ideal case, leave the ground state population unchanged. For this reason, these pulses are referred to as transitionless [43]. This phenomenon is by design, as it facilitates the transfer of coherences from the electronic states onto the emitted photons.
@@ III Stabilizer Generator Expectation Values
For a full quantum state tomography or detailed analysis of photon entanglement in a given $N$-photon quantum state, one might consider computing the $N$-photon density matrix directly. From here, protocol fidelities can be calculated [68]. However, in the microscopic simulations, this approach becomes computationally intractable, typically already for $N>3$, due to the exponential increase in space dimensionality and requirement for the explicit evaluation of N-th order photon correlation functions $\mathcal{G}^{(N)}$, also compare with [69] for details on $\mathcal{G}^{(1)}$ and $\mathcal{G}^{(2)}$. To keep the computational effort manageable, here we utilize stabilizer generator expectation values [62] as our primary metric for evaluating quantum correlations in the generated $N$-photon quantum states of interest. These values allow a localized investigation of entanglement by probing correlations in a segmented fashion, thereby significantly reducing computational demands. This is accomplished by evaluating correlation functions $\mathcal{G}^{(3)}$ only up to third-order, resulting in a tractable and efficient means to estimate the quantum entanglement present in the system without the need to reconstruct the full $N$-photon state. This effectively local approach not only simplifies the computational process but also delivers a detailed assessment of the entanglement properties. Stabilizer generator expectation values offer these capabilites to serve as benchmarks for these states because by definition, the graph states are composed of a superposition of stabilizer expressions [62, 70]. Unlike readily accessible metrics for generation protocol quality, such as rotation fidelities, stabilizer generator expectation values that are directly derived from the $N$-photon quantum state offer an accurate reflection of the system’s entanglement characteristics [8], even in lossy environments. To construct the stabilizer generator expectation values within the early and late basis specifically for a 1D linear cluster state, we explicitly calculate both second and third order photon correlation functions:
@@
where the time arguments $t_{i}$ are shifted by the time bin length $T$, as indicated by the $E$ (early) and $L$ (late) subsripts of the corresponding correlation functions (cf. Appendix B). The multi-time correlation functions are calculated by using the von-Neuman equation again on a modified density matrix using the quantum regression theorem [35, 71]. The operator $a^{(\dagger)}$ is the annihilation (creation) operator corresponding to the first ground- and excited state. We refer to the transition between these states as the cycling transition. We then reduce the correlation functions by integrating over all time dimensions such that
@@
The operator $\hat{\Phi}=\cos(\varphi)\hat{X}+i\sin(\varphi)\hat{Y}$ defines the phase $\varphi$ of the stabilizer generator expectation values which is linearly tunable by adjusting the phase $\vartheta$ of the rotation pulse. The numerator is normalized with
@@
The Pauli spin matrices $\hat{X}$,$\hat{Y}$ and $\hat{Z}$ are multiplied using the Kronecker product. These expressions are specific to 1D linear cluster states, grow linearly with the temporal photon entanglement [39], and can be used to calculate a lower bound for the generated degree of entanglement. In the ideal case, the stabilizer generator expectation values approach unity. The necessary configuration for the local bins to calculate any of these stabilizer generator expectation values adheres to the sequence
@@
where $t_{0}$ marks the commencement of the first bin associated with the expression, and its positioning is flexible within the time frame of the simulation to calculate later generator expectation values. For $t_{0}>0$, we shift the starting time of all of the integrals in Eq. 6. Further details are given in Appendix B.
@@
We now explore how rotation fidelities directly influence the quantum characteristics of the generated $N$-photon states, specified by the stabilizer generator expectation values. Given that it is always possible to identify a pulse phase optimizing both characteristics simultaneously, we select an arbitrary phase for our analysis, focusing solely on the magnitude of these properties. Initially, we introduce additional radiative loss exclusively to the transition states, previously set to zero, and adjust it from zero to $2.5\upmu$eV. The radiative loss rate for the cycling transition is kept constant. In Fig. 5, we observe that increased radiative losses in the transition state correspond to a direct decrease in the magnitudes of the stabilizers, exhibiting an almost linear relationship for the losses shown. Subsequently, we apply spin dephasing to the ground state coherences, again with rates ranging from zero to $2.5\upmu$eV. We note that these values lead to dephasing times significantly exceeding the typical achievable values for spin qubits [72] and provide a solid estimation for the lower bound for the achievable entanglement. In this scenario, we notice a rapid exponential decline in the stabilizer values even for relatively small dephasing rates. It is important to note that this detrimental effect of the ground state dephasing is not apparent in the temporal dynamics of the system, which for small loss values closely resembles the zero-dephasing behavior depicted in Fig. 2 b). It is only for losses exceeding $3\upmu$eV that a noticeable difference also becomes visible for the system dynamics.
@@
Figure 6: Local stabilizer generator expectation values and cluster-state length limitation. a) Magnitudes of the first 15 $\langle Z\Phi Z\rangle$ expectation values. The lossless case (blue line) is compared to either significant decay of the target and unwanted states (orange line) with $\gamma^{R}_{\text{T,U}}=0.5\cdot\gamma_{\text{Cyc}}\equiv 1$GHz, which corresponds to a lifetime of the transition states of $1$ns, or moderate spin dephasing of the ground states (green line) with $\gamma^{D}_{\text{T,U}}=0.0375\cdot\gamma_{\text{Cyc}}\equiv 75$MHz, which corresponds to a spin lifetime greater than $13$ns. b) Lower bound for the possible length of the linear cluster state according to Eq. 13, assuming all stabilizer expectation values are equal.
@@
It becomes evident that, although the rotation fidelity diminishes somewhat with increasing losses, the nonlinear decay of quantum correlations with increasing spin dephasing does only emerge in a detailed analysis of the quantum state properties through stabilizer generators. Note that we find that the selection of an optimal bin length is also influenced by the system’s loss parameters (not shown), with systems experiencing higher levels of loss potentially benefiting from shorter time bins, which helps to mitigate the adverse effects of dephasing over extended periods of time, optimizing the overall performance of the system under such conditions.